OpenStudy (anonymous):

Find the coordinate of point A of the line segment AB if the segment has an endpoint B(-2,10) and a midpoint at (3,-2). Am I supposed to use distance formula or slope to solve?

OpenStudy (anonymous):

slope

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

use 2 equations..

OpenStudy (anonymous):

10 = m(-2)+b -2=m(3)+b

OpenStudy (anonymous):

so your plugging in for both x and y 's

OpenStudy (anonymous):

then set the b's equal to each other, and your home free

OpenStudy (anonymous):

10+2m = -3m - 2

OpenStudy (anonymous):

5m = -12 m = -12/5

OpenStudy (anonymous):

thrn plug in that slope into either equation to solve for b

OpenStudy (anonymous):

10 = (-12/5)(-2) + b

OpenStudy (anonymous):

So ultimately what would the coordinate of point A be?